Consider a perfectly competitive market where Demand is described as Qd 100-2P. a. If the market price is 10, how many units are consumed in the market? What is the consumer surplus in the market? b. Suppose the market Supply is described as Qs 10 P. What is the equilibrium price in the market? Quantity? C. Suppose the market Supply is described as Qs 10+ P. What is the excess quantity supplied in the market at P demanded in the market at P 10? 40? What is the excess quantity d. Suppose the market Supply is described as Qs 10 P. Derive the excess supply curve in the market. (This question asks for a function that gives out the excess quantity supplied at every given price level.)
Answer #1

Answer a : P=$10

Q d = 100-2P= 100-2(10)= 80 units

Consumer surplus = 0.5(Maximum price- applicable price)( Maximum quantity- quantity)= 0.5(50-10)(80)=$1600

B Equilibrium

Quantity demanded= Quantity Supplied

100-2P= 10+P

100-10= 3P

90= 3P

P= $30

Equilibrium price =$30

Equilibrium quantity= 40 units

C: When price is $40 than quantity supplied = 10+P=10+40= 50 units

Excess supply units is 10.

When price is $10 than quantity demanded is 100-2P= 100-2( 10)=100-20=80 units

Excess quantity demanded as compared to Equilibrium is 40units.

Answer D : Excess supply curve shows the difference of 10 units when the price increases.

Excess supply curve = 10